# A flexible quantum data bus

### v1

@@ Appendix E Straight to diagonal routing of multiple data lines in a 2D cluster state
In this section we present the measurement pattern to change from parallel to diagonal data lines of a quantum data bus in a 2D cluster state. Applying the zipper-scheme to the last data line in a parallel transport together with the restoration property of the zipper-scheme enables to transport the next qubit along the seam closed by the zipper-scheme. The top part of Fig. 7 depicts the result after switching the direction of the last qubit of a horizontal quantum data bus to diagonal (blue) using the zipper-scheme. We transport the second-last qubit now by measuring along the seam (red line) produced by the zipper-scheme of the last qubit in the first step. The figure at the bottom of Fig. 7 shows the result after the both qubits of the quantum data bus switched the direction from horizontal to diagonal. Note that the length of the initial path for the zipper-scheme determines the number of qubits for the transfer, where a length of $2n$ enables to transfer $n$ qubits. Furthermore, we note that this approach inverts the order of data lines in the quantum data bus, similar as for parallel transport (see also Appendix C for more information).
Pauli X measurement
Pauli Z measurement
@@
Figure 7: Measurement pattern to change between parallel and diagonal information transport is shown. The qubit in the bottom of the parallel line stack has been transported first by with the zipper-scheme. The next qubit in the stack is transported by applying the zipper-scheme along the seam, produced by the first application of the zipper-scheme.
Appendix F ’V’ turn of multiple data lines in a 2D cluster state
In this section we demonstrate the measurement pattern to perform a ’V’ turn of parallel data lines. It basically boils down to applying the same measurement strategy as in the previous section of Appendix E, i.e. to change from straight to diagonal transport, just twice. In the top part of Fig. 8 we applied the zipper-scheme along the blue, dashed paths in order to propagate the bottom data line to the top. We use the closed seam (red) of the zipper-scheme, like for straight-to-diagonal transport, to propagate the next data line to the top. To propagate the qubits now from top to bottom again, we apply the previous measurement pattern for all data lines again, but now downwards in such a way that the resulting paths correspond to a ’V’. We note that we require one intermediate qubit per line at the turning point to enable the full functionality of the zipper-scheme. In the bottom part of Fig. 8 the intermediate result for a ’V’ turn of two parallel data lines to the top and one to the bottom is shown. In order to turn around $n$ data lines, a total number of $2(n\times n)$ is required.
Pauli X measurement
Pauli Z measurement
@@
Figure 8: Measurement strategy to perform a ’V’ turn on two parallel data lines. Top sketch illustrates the application of the zipper-scheme along the blue, dashed path to propagate the bottom data line to the top, and the next data line will be propagated to the top by applying the zipper-scheme along the red seam created the zipper-scheme. The bottom part sketches the ’V’ turn after both data lines have been brought up and the bottom line has been brought down. As a final step the two qubits in the middle have to be merged by Pauli X measurements.
Appendix G Fade-in and -out to a quantum data bus
We achieve a fade-in and -out to or from a quantum data bus by placing Pauli Z measurements at appropriate positions. In Fig. 9 we show a fade-out of a three qubit data line bundle where the yellow measurements correspond to Pauli Z measurements for cutting out the data lines, which are merged subsequently by the orange Pauli X measurements.
Pauli X measurement
Pauli Z measurement
@@ Appendix G Fade-in and -out to a quantum data bus
Figure 9: Pauli Z measurements are depicted by the yellow badges, and those cut out the measurement path. By measuring the qubits along the orange lines in the Pauli X basis, we achieve the fade-out of the three data lines.
Appendix H Zipper-scheme for one-dimensional cluster state and GHZ state extraction
In this section, we show how to apply the zipper-scheme to extract one-dimensional cluster states and GHZ states from a two dimensional cluster state. We point out that by using insights from Ref. [51] together with our strategy we generate GHZ states. The simplest way to extract a linear cluster state from a two dimensional cluster state is by applying the zipper-scheme and not measuring qubits on the path. In the first part from the left of Fig. 10 we show how to extract a GHZ state or a three qubit linear cluster state from a 2D cluster state by applying the zipper-scheme along the purple measurement path. However, we note that this comes with an increased number of qubits we have to measure in the Pauli Z basis (yellow qubit along the seam), because (green) edges to the target state remain in addition to the (red) edges due to the zipper-scheme. Overall, the additional Pauli Z measurements of qubits reduce the connectivity in the remaining resource state.
Pauli X measurement
Pauli Z measurement
@@ Appendix H Zipper-scheme for one-dimensional cluster state and GHZ state extraction
Figure 10: Linear cluster or GHZ extraction: In the first sub-figure we apply the zipper-scheme along a (purple, dashed) staircase-shaped path to extract a one-dimensional cluster or GHZ state (purple qubits) from the 2D cluster state by keeping intermediate qubits unmeasured, and measuring additional qubits (yellow) within the Pauli Z basis. In the middle we demonstrate that the remaining cluster state, due to the zipper-scheme, enables to extract two one-dimensional cluster or GHZ state adjacent to each other. In the right part an additional one-dimensional cluster or GHZ state (turquoise) is extracted orthogonal to two established states.
H.1 Crossing of 1D cluster states and GHZ states
@@
In summary, using the zipper-scheme to distill a linear cluster or GHZ state from the 2D cluster works, but demands a measurement for each additional qubit we add to the Bell pair. Additionally, the generation of GHZ states [51] with more than three qubits from the linear cluster state requires approximately half of the qubits to measured.

### v3

@@ Appendix D Straight to diagonal routing of multiple data lines in a 2D cluster state
In this section we present the measurement pattern to change from parallel to diagonal data lines of a quantum data bus in a 2D cluster state. Applying the zipper scheme to the last data line in a parallel transport together with the restoration property of the zipper scheme enables to transport the next qubit along the seam closed by the zipper scheme. The top part of Fig. 7 depicts the result after switching the direction of the last qubit of a horizontal quantum data bus to diagonal (blue) using the zipper scheme. We transport the second-last qubit now by measuring along the seam (red line) produced by the zipper scheme of the last qubit in the first step. The figure at the bottom of Fig. 7 shows the result after the both qubits of the quantum data bus switched the direction from horizontal to diagonal. Note that the length of the initial path for the zipper scheme determines the number of qubits for the transfer, where a length of $2n$ enables to transfer $n$ qubits. Furthermore, we note that this approach inverts the order of data lines in the quantum data bus, similar as for parallel transport (see also Appendix C for more information).
@@
Figure 7: Measurement pattern to change between parallel and diagonal information transport is shown. The qubit in the bottom of the parallel line stack has been transported first by with the zipper scheme. The next qubit in the stack is transported by applying the zipper scheme along the seam, produced by the first application of the zipper scheme.
@@
Figure 8: Measurement strategy to perform a 'V'-turn on two parallel data lines. Top sketch illustrates the application of the zipper scheme along the blue, dashed path to propagate the bottom data line to the top, and the next data line will be propagated to the top by applying the zipper scheme along the red seam created the zipper scheme. The bottom part sketches the 'V'-turn after both data lines have been brought up and the bottom line has been brought down. As a final step the two qubits in the middle have to be merged by Pauli $X$ measurements.
Appendix E 'V'-turn of multiple data lines in a 2D cluster state
In this section we demonstrate the measurement pattern to perform a 'V'-turn of parallel data lines. It basically boils down to applying the same measurement strategy as in the previous section of Appendix D, i.e. to change from straight to diagonal transport, just twice. In the top part of Fig. 8 we applied the zipper scheme along the blue, dashed paths in order to propagate the bottom data line to the top. We use the closed seam (red) of the zipper scheme, like for straight-to-diagonal transport, to propagate the next data line to the top. To propagate the qubits now from top to bottom again, we apply the previous measurement pattern for all data lines again, but now downwards in such a way that the resulting paths correspond to a 'V'. We note that we require one intermediate qubit per line at the turning point to enable the full functionality of the zipper scheme. In the bottom part of Fig. 8 the intermediate result for a 'V'-turn of two parallel data lines to the top and one to the bottom is shown. To turn around $n$ data lines, a total number of $2(n\times n)$ is required.
@@
Figure 9: Pauli $Z$ measurements are depicted by the yellow badges, and those cut out the measurement path. By measuring the qubits along the orange lines in the Pauli $X$ basis, we achieve the splitting of the three data lines.
@@
Figure 10: Linear cluster or GHZ extraction: In the first sub-figure we apply the zipper scheme along a (purple, dashed) staircase-shaped path to extract a one-dimensional cluster or GHZ state (purple qubits) from the 2D cluster state by keeping intermediate qubits unmeasured, and measuring additional qubits (yellow) within the Pauli $Z$ basis. In the middle we demonstrate that the remaining cluster state, due to the zipper scheme, enables to extract two one-dimensional cluster or GHZ state adjacent to each other. In the right part an additional one-dimensional cluster or GHZ state (turquoise) is extracted orthogonal to two established states.
Appendix F Merging and splitting of data lines of the data bus
We achieve a merging and splitting of the data lines that constitute the quantum data bus by placing Pauli $Z$ measurements at appropriate positions. In Fig. 9 we show a splitting of a three qubit data line bundle where the yellow measurements correspond to Pauli $Z$ measurements for cutting out the data lines, which are merged subsequently by the orange Pauli $X$ measurements.
Appendix G Zipper scheme for one-dimensional cluster state and GHZ state extraction
In this section, we show how to apply the zipper scheme to extract one-dimensional cluster states and GHZ states from a two dimensional cluster state. We point out that by using insights from Ref. [52] together with our strategy we generate GHZ states. The simplest way to extract a linear cluster state from a two dimensional cluster state is by applying the zipper scheme and not measuring qubits on the path. In the first part from the left of Fig. 10 we show how to extract a GHZ state or a three qubit linear cluster state from a 2D cluster state by applying the zipper scheme along the purple measurement path. However, we note that this comes with an increased number of qubits we have to measure in the Pauli $Z$ basis (yellow qubit along the seam), because (green) edges to the target state remain in addition to the (red) edges due to the zipper scheme. Overall, the additional Pauli $Z$ measurements of qubits reduce the connectivity in the remaining resource state.
G.1 Crossing of 1D cluster states and GHZ states
@@
In summary, using the zipper scheme to distill a linear cluster or GHZ state from the 2D cluster works, but demands a measurement for each additional qubit we add to the linear cluster state. Additionally, the generation of GHZ states [52] with more than three qubits from the linear cluster state requires approximately half of the qubits to measured.